# Fractional Dynamics in Galactic Nuclei Non-Local Transport, Transient Phenomena and the Nullification of the Schwarzschild Barrier

### v1

@@ 6.3 Fractional Dynamics: Non-Local Barrier Jumping
The fractional framework provides a fundamentally different result due to the non-local nature of the transport operators. This distinction resolves the bottleneck suggested by the local models. We demonstrate this mathematically by analyzing the structure of the fractional derivatives.
@@
6.4 The Nullification of the Barrier
The consequence of this non-local transport is profound. The bottleneck predicted by local models does not inhibit the flux in a non-local framework. The flux into the loss cone is maintained by direct jumps from the region $J>J_{\text{SB}}$. As shown in Figure 3, the fractional flux remains unsuppressed ($\mathcal{F}_{\text{frac}}=1.0$), completely overcoming the 88% reduction predicted by the local model. Therefore, the proposed Schwarzschild Barrier, while representing a real physical phenomenon of local quenching via AI, does not constitute an actual barrier to RR-driven transport.
This theoretical result is consistent with observations from direct-summation N-body simulations using integrators, as detailed in the next section.
Since RR can bypass the quenched region via non-local jumps, and two-body relaxation is inherently unaffected by precession, we conclude that the Schwarzschild Barrier does not impose any significant inhibition on the production of EMRIs (P. Amaro-Seoane, 2018, 2020; P. Amaro-Seoane et al., 2007). This result resolves a major theoretical bottleneck in EMRI formation theory.
@@ 7 Numerical Evidence from N-body Simulations
To validate the theoretical prediction that the Schwarzschild Barrier (SB) is ineffective against non-local transport, we analyze the results of direct-summation $N$-body simulations incorporating relativistic corrections, following the pioneering work of (G. Kupi et al., 2006). The full details (method, procedure and setup) of the numerical methods, including the post-Newtonian (PN) terms, are available in P. Brem et al. (2014).
@@ 7.1 Simulation Setup and Methodology
The simulations model a central SMBH with a mass of $M_{\bullet}=10^{6}M_{\odot}$ surrounded by $N=400$ stellar-mass black holes (BHs), each with a mass of $m_{\star}=50M_{\odot}$. The stellar BHs are initialized with a flat distribution in semi-major axis $a$ and $e^{2}$, with $a$ ranging from $0.1\rm{mpc}$ to $10\rm{mpc}$.
This range of semi-major axes is crucial. For a $10^{6}M_{\odot}$ SMBH, the radius of influence $r_{\rm inf}$ is typically on the order of parsecs. The simulated domain (milliparsecs) is therefore deep within the SMBH’s sphere of influence. In this region, the enclosed stellar mass $M_{\text{cusp}}(r)$ is much smaller than $M_{\bullet}$. According to Eq. (10), this ensures that Resonant Relaxation (RR) is the dominant transport mechanism, significantly outweighing two-body relaxation. Simultaneously, this proximity is where relativistic precession becomes rapid, defining the location of the SB. These simulations thus provide an ideal testbed for assessing the impact of the Schwarzschild Barrier in the regime where it is hypothesized to be most restrictive.
The figures presented here are generated from a statistical sample of $N_{\rm simul}=300$ such simulations. We generate a 2D histogram by sampling the orbital elements (semi-major axis $a$ and eccentricity $e$) of all $N=400$ particles across all 300 simulations throughout their evolution. The orbital elements are binned into a $100\times 100$ grid in $\log_{10}(a)$ versus $\log_{10}(1-e)$ phase space. The count in each bin is then normalized by the total number of samples and the bin area to create a presence density map, $\log_{10}P(a,1-e)$, which represents the time-averaged probability of finding a particle in a given region of phase space.
@@ 7.2 Results and Interpretation
Figure 4: Time-averaged presence density of stellar-mass black holes in the $a$ vs. 1e1-e phase space, generated from 300 NN-body simulations of 400 particles each, focusing on the critical milliparsec scales where RR dominates and precession is strong. Left panel: The purely Newtonian case (no PN terms). Particles are driven to high eccentricities (low 1e1-e) by relaxation processes, filling the phase space. Right panel: The relativistic (PN) case without spin. The solid blue line marks the theoretical location of the Schwarzschild Barrier ($\tau_{\rm SS}\approx\tau_{\rm coh}$), and the dashed blue line marks the gravitational wave capture limit ($\tau_{\rm GW}<\tau_{\rm rlx}$). The particle density is high on both sides of the SB (solid line), with no visible depletion or barrier to transport. The distribution is truncated only by the GW capture line, confirming that the SB does not inhibit particle flow into the loss cone.
@@
It is visually evident from the right panel that the Schwarzschild Barrier does not function as an actual barrier to transport. The presence density (dark red shading) is high on both sides of the solid blue line. This demonstrates that particles continuously and efficiently cross this theoretical boundary from the region where RR is locally efficient (right) into the region where it is locally quenched by precession (left). The fact that this efficient transport occurs in a regime where RR is the dominant mechanism emphasizes that the SB is porous to RR-driven flux. The particle distribution is not truncated or depleted at the SB. Instead, the density extends all the way to the GW capture limit (dashed line), where particles are removed from the system. This provides direct numerical evidence supporting the non-local barrier jumping mechanism derived mathematically from the fractional formalism (Section 6).
@@ 8 Conclusions
We have established the necessity of a fractional dynamics framework for accurately modeling non-local transport in stellar nuclei. We demonstrated that superdiffusive processes are Lévy flights characterized by infinite variance and heavy-tailed jump distributions. This physical reality contravenes the assumptions of the standard local Fokker-Planck equation, which is mathematically restricted to Gaussian processes governed by the Central Limit Theorem. By applying the space-Fractional Fokker-Planck Equation (space-FFPE), which employs non-local integro-differential operators (Eq. 8), we correctly model the process according to the Generalized Central Limit Theorem. This fractional framework resolves significant discrepancies between local theory and astrophysical observations. We showed that local models predict an unphysical exponential delay in transient loss cone refilling ($\Gamma(t)\propto\exp(-C/t)$), failing to explain high TDE rates in post-starburst galaxies. The fractional model, by contrast, predicts an immediate linear flux ($\Gamma(t)\propto t$) (Section 5). Furthermore, we proved mathematically that the non-local integral structure of the fractional operators permits “barrier jumping” (Section 6). This mechanism nullifies the postulated Schwarzschild Barrier as a transport bottleneck. This theoretical finding is confirmed by direct $N$-body simulations incorporating relativistic effects (Section 7), which explicitly show that the particle presence density is continuous across the theoretical barrier and is truncated only by the gravitational-wave capture limit (Figure 4). Local models, bound by differential operators, misinterpret this local quenching of diffusion as a severe barrier to EMRI production (B. Bar-Or & T. Alexander, 2016). Our combined theoretical and numerical findings demonstrate that this barrier does not inhibit EMRI rates, resolving a major theoretical bottleneck.
@@
The implementation of this fractional framework into existing stellar dynamics codes presents a significant but necessary numerical challenge. Standard Fokker-Planck solvers are typically built on finite-difference schemes, such as the Chang-Cooper method, which discretize a local differential operator ($\mathcal{L}_{\text{FP}}$). This results in sparse (e.g., tridiagonal) matrices that are computationally efficient to solve. The fractional operators, however, are non-local integro-differential operators. Their discretization, for instance, can be achieved using a fractional finite difference method. This method approximates the Riesz operator $\mathbb{D}_{x}^{\alpha}f(x)$ by $h^{-\alpha}\Delta^{\alpha}f(x)$, where $\Delta^{\alpha}$ is a fractional centered-difference operator derived from the Grünwald-Letnikov series $\Delta^{\alpha}f(x)=\sum_{j=0}^{\infty}\binom{\alpha}{j}(-1)^{j}f(x-jh)$ (M. M. Meerschaert & A. Sikorskii, 2011). This discretization, which couples $f(x)$ to all points f(xjh)f(x-jh), is what leads to the dense matrix system and its $O(N^{2})$ (or even worse) scaling. However, stable implicit finite difference codes can be constructed, for example by shifting the index of the numerical operator, which is essential for robust numerical solutions (M. M. Meerschaert & A. Sikorskii, 2011). Despite this computational cost, this transition is necessary. The results of this work demonstrate that local approximations are not merely inaccurate; they are qualitatively incorrect, predicting spurious bottlenecks and time delays that are artifacts of the model rather than the physics. Adopting these non-local numerical methods is therefore imperative for producing physically correct simulations of event rates in galactic nuclei.

### v2

@@ 7.3 Fractional Dynamics: Non-Local Barrier Jumping
The fractional framework provides a fundamentally different result due to the non-local nature of the transport operators. This distinction potentially resolves the bottleneck suggested by the local models. We demonstrate this mathematically by analyzing the structure of the fractional derivatives.
@@
7.4 The Nullification of the Barrier and Reinterpretation of Prior Results
The consequence of this non-local transport is profound. The bottleneck predicted by local models does not inhibit the flux in a non-local framework. The flux into the loss cone is maintained by direct jumps from the region $J>J_{\text{SB}}$. As shown in Figure 3, the fractional flux remains unsuppressed ($\mathcal{F}_{\text{frac}}=1.0$), completely overcoming the 88% reduction predicted by the local model. Therefore, the proposed Schwarzschild Barrier, while representing a real physical phenomenon of local quenching via AI, does not constitute an actual barrier to RR-driven transport in this framework.
The apparent discrepancy between this theoretical result and the numerical observations of a barrier in D. Merritt et al. (2011) (MAMW11) and A. S. Hamers et al. (2014) warrants careful consideration. We argue this discrepancy stems primarily from the methodologies used to analyze the simulations, which inherently assume local transport.
A. S. Hamers et al. (2014) analyzed the dynamics by measuring the diffusion coefficients (DCs), $D(L)=\langle(\Delta L)^{2}\rangle/\Delta t$. They correctly observed a sharp suppression (a ”knee”) in $D(L)$ near the SB. However, if RR is a Lévy flight, the second moment $\langle(\Delta L)^{2}\rangle$ formally diverges (Section 2). In any finite simulation, the measured moments are necessarily finite because the power-law tail is truncated. The measured $D(L)$ is therefore an effective diffusion coefficient, dominated by the frequent small steps and under-representing the rare large jumps (R. Metzler & J. Klafter, 2000).
The suppression measured by A. S. Hamers et al. (2014) reflects the quenching of local transport (small steps) due to Adiabatic Invariance. The crucial non-local jumps, which bypass the barrier, are too infrequent to significantly contribute to their measurement of $D(L)$ over the analyzed timescales. Consequently, analyzing the system via local diffusion coefficients leads to the interpretation that the overall transport is suppressed, even when non-local flux persists.
In the case of MAMW11, the barrier was identified visually by the ”bouncing” of trajectories. This is expected even in the fractional framework, as most steps are small and affected by local quenching. Furthermore, MAMW11 utilized very small N ($N=50$), enhancing TBR relative to RR. MAMW11 concluded that the few captures they observed were likely driven by TBR, not RR.
Since RR can bypass the quenched region via non-local jumps, and two-body relaxation is inherently unaffected by precession, we conclude that the Schwarzschild Barrier may not impose the significant inhibition previously assumed on the production of EMRIs (P. Amaro-Seoane, 2018, 2020; P. Amaro-Seoane et al., 2007). This result addresses a major theoretical bottleneck in EMRI formation theory.
@@ 8 Numerical Evidence from N-body Simulations
To test the theoretical prediction that the Schwarzschild Barrier (SB) is ineffective against non-local transport, we analyze the results of direct-summation $N$-body simulations incorporating relativistic corrections, following the pioneering work of (G. Kupi et al., 2006). The full details of the numerical methods, including the post-Newtonian (PN) terms, are available in P. Brem et al. (2014).
@@ 8.1 Simulation Setup and Methodology
The simulations model a central SMBH with a mass of $M_{\bullet}=10^{6}M_{\odot}$ surrounded by $N=400$ stellar-mass black holes (BHs), each with a mass of $m_{\star}=50M_{\odot}$. The stellar BHs are initialized with a flat distribution in semi-major axis $a$ and $e^{2}$ (a thermal distribution), with $a$ ranging from $0.1\rm{mpc}$ to $10\rm{mpc}$.
These simulations are intended as a proof of concept and possess significant limitations. The particle number ($N=400$) is very small compared to realistic galactic nuclei, and the individual stellar mass ($m_{\star}=50M_{\odot}$) is large. This significantly enhances the efficiency of two-body relaxation (TBR) relative to RR. Furthermore, the initialization with a thermal eccentricity distribution means the system starts near equilibrium, rather than demonstrating the dynamic process of diffusing across the barrier from low eccentricities.
This range of semi-major axes is crucial. For a $10^{6}M_{\odot}$ SMBH, the radius of influence $r_{\rm inf}$ is typically on the order of parsecs. The simulated domain (milliparsecs) is therefore deep within the SMBH’s sphere of influence. In this region, the enclosed stellar mass $M_{\text{cusp}}(r)$ is much smaller than $M_{\bullet}$. According to Eq. (16), this ensures that Resonant Relaxation (RR) is the dominant transport mechanism, significantly outweighing two-body relaxation, provided N is sufficiently large and $m_{\star}$ sufficiently small. Simultaneously, this proximity is where relativistic precession becomes rapid, defining the location of the SB.
The figures presented here are generated from a statistical sample of $N_{\rm simul}=300$ such simulations. The simulations are run for a duration sufficient to establish a quasi-steady state distribution. We generate a 2D histogram by sampling the orbital elements (semi-major axis $a$ and eccentricity $e$) of all $N=400$ particles across all 300 simulations throughout their evolution. The orbital elements are binned into a $100\times 100$ grid in $\log_{10}(a)$ versus $\log_{10}(1-e)$ phase space. The count in each bin is then normalized by the total number of samples and the bin area to create a presence density map, $\log_{10}P(a,1-e)$, which represents the time-averaged probability of finding a particle in a given region of phase space.
@@ 8.2 Results and Interpretation
Figure 4: Time-averaged presence density of stellar-mass black holes in the $a$ vs. 1e1-e phase space, generated from 300 NN-body simulations of 400 particles each, focusing on the critical milliparsec scales. Left panel: The purely Newtonian case (no PN terms). Particles are driven to high eccentricities (low 1e1-e) by relaxation processes, filling the phase space. Right panel: The relativistic (PN) case without spin. The solid blue line marks the theoretical location of the Schwarzschild Barrier ($\tau_{\rm SS}\approx\tau_{\rm coh}$), and the dashed blue line marks the gravitational wave capture limit ($\tau_{\rm GW}<\tau_{\rm rlx}$). The particle density is high on both sides of the SB (solid line), with no visible depletion or barrier to transport in this TBR-dominated regime.
@@
It is visually evident from the right panel that the Schwarzschild Barrier does not function as an actual barrier to transport in these simulations. The presence density (dark red shading) is high on both sides of the solid blue line.
However, the interpretation of this result requires careful consideration of the relaxation regimes, as analyzed by A. S. Hamers et al. (2014). They demonstrated that the impact of the SB depends on the relative strength of TBR versus the suppressed RR (which they termed Anomalous Relaxation, AR) below the barrier. They defined a critical radius, $a_{\mathrm{AR,max}}$ above which TBR is strong enough to dominate transport across the SB. For the parameters used here ($N=400,m_{\star}=50M_{\odot},M_{\bullet}=10^{6}M_{\odot}$), we estimate $a_{\mathrm{AR,max}}\approx 1.4$ mpc.
Therefore, for the majority of the simulated phase space ($a>1.4$ mpc), the efficient transport observed across the SB is likely dominated by TBR, which is inherently unaffected by relativistic precession. The simulations are in a regime where enhanced TBR (due to low N and high $m_{\star}$) masks the effects of the SB on RR. Consequently, while these results show efficient transport, they cannot be used to conclusively validate the hypothesis that RR bypasses the barrier via non-local jumps.
The particle distribution is not truncated or depleted at the SB. Instead, the density extends all the way to the GW capture limit (dashed line), where particles are removed from the system. This provides numerical evidence of efficient transport across the SB.
To robustly validate the fractional diffusion model, a comprehensive numerical investigation is required. This investigation must include simulations with significantly larger $N$ and smaller $m_{\star}$ to ensure the dominance of RR and reduce the impact of TBR (ensuring $a\ll a_{\mathrm{AR,max}}$), exploration of different initial conditions (e.g., starting from quasi-circular orbits far from the barrier), and, critically, a direct measurement of the angular momentum jump probability distribution $P(\Delta J)$ from the simulations. This measurement is essential to confirm the presence of the heavy tails predicted by the Lévy flight hypothesis (Section 2). Such an extensive study requires significant computational resources and specialized analysis techniques; it is beyond the scope of this theoretical paper and will be presented in a separate, dedicated work.
@@ 9 Conclusions
We have investigated the application of a fractional dynamics framework for modeling transport in stellar nuclei. We provided a formal justification (Section 2) that Resonant Relaxation (RR) must be characterized as a Lévy flight due to the infinite variance inherent in the gravitational torques. We utilized established results from statistical physics (e.g., M. M. Meerschaert & A. Sikorskii, 2011) that superdiffusive processes (Lévy flights) are characterized by infinite variance and heavy-tailed jump distributions. This physical reality contravenes the assumptions of the standard local Fokker-Planck equation, which is mathematically restricted to Gaussian processes governed by the Central Limit Theorem. By applying the space-Fractional Fokker-Planck Equation (space-FFPE), which employs non-local integro-differential operators (Eq. 14), we provide a model consistent with the Generalized Central Limit Theorem.
This fractional framework offers potential resolutions to significant discrepancies between local theory and astrophysical observations. We showed that local models predict a significant exponential delay in transient loss cone refilling ($\Gamma(t)\propto\exp(-C/t)$), failing to explain high TDE rates in post-starburst galaxies. The fractional model, by contrast, predicts an immediate linear flux ($\Gamma(t)\propto t$) (Section 6).
Furthermore, we demonstrated analytically that the non-local integral structure of the fractional operators permits “barrier jumping” (Section 7). This mechanism bypasses the postulated Schwarzschild Barrier as a transport bottleneck. We presented preliminary, proof-of-concept direct $N$-body simulations incorporating relativistic effects (Section 8), which indicate that the particle presence density is continuous across the theoretical barrier. However, we analyzed the relaxation regimes following A. S. Hamers et al. (2014) and concluded that this observed efficiency is likely due to enhanced TBR (due to small $N$ and large $m_{\star}$) rather than non-local RR. Further investigation with large-N simulations, including direct measurement of the jump statistics, is required to validate the non-local RR hypothesis and will be presented in a forthcoming work.
Local models, bound by differential operators, misinterpret this local quenching of diffusion as a severe barrier to EMRI production (B. Bar-Or & T. Alexander, 2016). Our theoretical findings strongly suggest that if RR is indeed a Lévy flight, this barrier may not inhibit EMRI rates as severely as previously thought, potentially resolving a major theoretical bottleneck.
@@
The implementation of this fractional framework into existing stellar dynamics codes presents a significant but necessary numerical challenge. Standard Fokker-Planck solvers are typically built on finite-difference schemes, such as the Chang-Cooper method, which discretize a local differential operator ($\mathcal{L}_{\text{FP}}$). This results in sparse (e.g., tridiagonal) matrices that are computationally efficient to solve. The fractional operators, however, are non-local integro-differential operators. Their discretization, for instance, can be achieved using a fractional finite difference method. This method approximates the Riesz operator $\mathbb{D}_{x}^{\alpha}f(x)$ by $h^{-\alpha}\Delta^{\alpha}f(x)$, where $\Delta^{\alpha}$ is a fractional centered-difference operator derived from the Grünwald-Letnikov series $\Delta^{\alpha}f(x)=\sum_{j=0}^{\infty}\binom{\alpha}{j}(-1)^{j}f(x-jh)$ (M. M. Meerschaert & A. Sikorskii, 2011). This discretization, which couples $f(x)$ to all points f(xjh)f(x-jh), is what leads to the dense matrix system and its $O(N^{2})$ (or even worse) scaling. However, stable implicit finite difference codes can be constructed, for example by shifting the index of the numerical operator, which is essential for robust numerical solutions (M. M. Meerschaert & A. Sikorskii, 2011). Despite this computational cost, this transition appears warranted. The results of this work demonstrate that local approximations may be qualitatively incorrect, potentially predicting spurious bottlenecks and time delays that are artifacts of the model rather than the physics. Adopting these non-local numerical methods is therefore imperative for producing physically accurate simulations of event rates in galactic nuclei.